The lifespan of housecats is normally distributed with a mean of 15.1 years and a standard deviation of 1.75 years

a. Find the age of a housecat that separates the oldest​ 8% from the youngest​ 92%

b. describe the location and shaded region under the normal curve
Answer #1

solution Given that el=isol years. 0 = 1.25 Years. Using normal distribution table Plz> Z) = 8% = 1-pcz<3)0.08 -P( ZCZ) = 1-0